OpenStudy (anonymous):

what is the answer to this? y=3x x 3x+x+1

OpenStudy (anonymous):

\[\gamma=3\chi^{2}+\chi+1\]

OpenStudy (anonymous):

How do u graph that?

OpenStudy (anonymous):

please help me out with it!!!!

OpenStudy (amistre64):

Is the equation: y= 3x (3x +x +1)?

OpenStudy (amistre64):

3x^2 +x +1 To vind the vertex, it is easier to see it in the form: y = m(x+h)^2 +k (h,k) would be the vertex.

OpenStudy (anonymous):

vertex point (r,k) ax^2+bx+c=0 then r = -b/2a k=f(r)

OpenStudy (amistre64):

In calculus you take its derivative and make it equal to 0: 6x + 1 = 0 when x= -1/6 3(-1/6)^2 + (-1/6) +1 =y 3/36 - 1/6 + 6/6 = y 3/36 - 6/36 + 36/36 = 33/36 = 11/12 The vertex should be at ((-1/6),(11/12)) if I did everything right :)

OpenStudy (anonymous):

thank you so much!